# Studying Ionospheric Phase Structure Functions Using Wide-Band uGMRT (Band-4) Interferometric Data

### v1

@@ Abstract
Interferometric observations of the low-frequency radio sky ($<$ 1 GHz) are largely limited by systematic effects introduced by the ionosphere. In this study, we use a ten-hour nighttime observation of the bright radio source 3C48, carried out with the upgraded Giant Metrewave Radio Telescope (uGMRT), to examine how ionospheric conditions affect radio signals. We focus on measuring phase fluctuations caused by the ionosphere and analyse how these variations change with antenna separation using the spatial phase structure function. Our results, based on data from three sub-bands between 575 and 725 MHz, show a clear power-law trend consistent with turbulent behaviour. We introduce the diffractive scale as a scalar measure of ionospheric conditions, which could guide future calibration strategies. Furthermore, we find that the turbulence is anisotropic, with properties that vary by direction. The smallest diffractive scales do not always align with Earth’s magnetic field, suggesting the influence of medium-scale Travelling Ionospheric Disturbances (MSTIDs). The contribution of these wave-like MSTIDs to the structure function follows a power-law with a slope close to 2.0, indicating their dominant role in shaping the observed anisotropic phase fluctuations.
@@ 2 Observation and Processing
We conducted 10 hours of nighttime observations of the bright quasar 3C48 using the upgraded GMRT (uGMRT) at Band-4 (project code: $47\_003$  [^1] https://naps.ncra.tifr.res.in/goa/data/search) in November 2024. The data span a 200 MHz bandwidth (550–750 MHz) with a central frequency of 650 MHz, using 8192 channels at 24.41 kHz resolution and an integration time of 2.68 seconds. For ease of computation and to avoid frequency ranges significantly affected by radio frequency interference (RFI), we selected a subset of three relatively clean sub-bands: 575–600 MHz, 600–625 MHz, and 700–725 MHz. We found overall good agreement among the analysed bands. More details about the observations are summarized in Table 1.
@@ 2.1 Calibration and Phase Processing
To isolate ionospheric fluctuations from slowly varying instrumental phases, we applied a one-hour boxcar smoothing filter to the unwrapped phases, treating the residual as ionospheric variability. Although absolute instrumental phases could not be fully separated, this approach enabled analysis of short-timescale TEC variations ($\leq$ 1 hour). Residual phase variations post-continuum subtraction were typically within ±$\pm 0.5$ radians for central square antennas and up to ±$\pm 5.0$ radians for arm antennas (Banerjee et al., 2025). A notable increase in residual scatter between 2-4 hours suggests enhanced ionospheric activity or unflagged RFI, consistent with scintillation reports in the uGMRT observation log. These fluctuations were significantly stronger than those in adjacent time steps, leading us to consider the data unreliable and flag this time interval (Banerjee et al., 2025).
3 Ionospheric Phase Structure Function
@@
Figure 1: Differential TEC ($\delta{\rm TEC}$) for the RR polarization of uGMRT baselines relative to the reference antenna ‘C06’ as a function of time. The colour bar indicates baseline length. (Top row) Baselines within the central square. (Bottom row) Extended antenna arm baselines.
@@ 4 Summary and Conclusions
To further assess anisotropy in the phase fluctuations, we compared the phase structure function along baselines that are approximately aligned with the Earth’s magnetic field (0–20) and those nearly perpendicular to it (70–90). Our results show that the diffractive scale is significantly larger in the aligned direction than in the perpendicular one (Figure 5). This indicates that the ionospheric irregularities in our observation are not aligned with the magnetic field lines.
@@
It is important to note that this study used only nighttime observations taken during the winter season. Therefore, our results do not account for diurnal or seasonal variations. Overall, our results demonstrate the capability of the uGMRT for ionospheric studies at mid-latitudes close to the geomagnetic equator (Opio et al., 2015).
@@ Appendix A Sub-Band Variations in Ionospheric Phase Structure Function (Band-4)
In this appendix, we present additional structure function analyses at different sub-bands within the uGMRT Band-4 frequency range to assess the frequency dependence of ionospheric phase fluctuations. Figures 7, 8, and 9 illustrate the results for two representative frequencies: 612.5 MHz and 712.5 MHz. These include the one-dimensional phase structure function with power-law fits, two-dimensional anisotropic structure functions highlighting directional differences in turbulence, and binned phase variance plots as a function of baseline orientation relative to the projected Earth’s magnetic field. Together, these plots provide insights into how the ionospheric irregularities and anisotropic characteristics evolve across Band-4 frequencies.

### v2

@@ Abstract
Interferometric observations of the low-frequency radio sky (<< 1 GHz) are largely limited by systematic effects introduced by the ionosphere. In this study, we use a ten-hour nighttime observation of the bright radio source 3C48, carried out with the upgraded Giant Metrewave Radio Telescope (uGMRT), to examine how ionospheric conditions affect radio signals. We focus on measuring phase fluctuations caused by the ionosphere and analyse how these variations change with antenna separation using the spatial phase structure function. Our results, based on data from three sub-bands between 575 and 725 MHz, show a clear power-law trend consistent with turbulent behaviour. We introduce the diffractive scale as a scalar measure of ionospheric conditions, which could guide future calibration strategies. Furthermore, we find that the turbulence is anisotropic, with properties that vary by direction. The smallest diffractive scales do not always align with Earth’s magnetic field, suggesting the influence of medium-scale Travelling Ionospheric Disturbances (MSTIDs). The contribution of these wave-like MSTIDs to the structure function follows a power-law with a slope close to 2.0, indicating their dominant role in shaping the observed anisotropic phase fluctuations.
@@ 2 Observation and Processing
We conducted 10 hours of nighttime observations of the bright quasar 3C48 using the upgraded GMRT (uGMRT) at Band-4 (project code: 47_00347\_003  [^1] https://naps.ncra.tifr.res.in/goa/data/search) in November 2024. The data span a 200 MHz bandwidth (550–750 MHz) with a central frequency of 650 MHz, using 8192 channels at 24.41 kHz resolution and an integration time of 2.68 seconds. For ease of computation and to avoid frequency ranges significantly affected by radio frequency interference (RFI), we selected a subset of three relatively clean sub-bands: 575–600 MHz, 600–625 MHz, and 700–725 MHz. We found overall good agreement among the analysed bands. More details about the observations are summarized in Table 1.
@@ 2.1 Calibration and Phase Processing
To isolate ionospheric fluctuations from slowly varying instrumental phases, we applied a one-hour boxcar smoothing filter to the unwrapped phases, treating the residual as ionospheric variability. Although absolute instrumental phases could not be fully separated, this approach enabled analysis of short-timescale TEC variations ($\leq$ 1 hour). Residual phase variations post-continuum subtraction were typically within $\pm 0.5$ radians for central square antennas and up to $\pm 5.0$ radians for arm antennas (Banerjee et al., 2025). A notable increase in residual scatter between 2-4 hours suggests enhanced ionospheric activity or unflagged RFI, consistent with scintillation reports in the uGMRT observation log. These fluctuations were significantly stronger than those in adjacent time steps, leading us to consider the data unreliable and flag 45 minutes of data particularly during the interval from 20:15 to 20:55 IST (Banerjee et al., 2025).
3 Results: Ionospheric Phase Structure Function
@@
Figure 1: Differential TEC ($\delta{\rm TEC}$) for the $R$ polarization of uGMRT baselines relative to the reference antenna ‘C06’ as a function of time. The colour bar indicates baseline length. (Top row) Baselines within the central square. (Bottom row) Extended antenna arm baselines.
@@ 3.2 Anisotropy in the ionospheric structures
Often, the phase structure function at short baselines hits a noise floor (Mevius et al., 2016). This is mainly due to second-order contributions from calibration phases (which are not taken into account) and other residual systematic errors. Although this effect is not very explicit in our uGMRT results (Figure 5), we do detect hints of flattening at shorter baselines (<1<1 km). The power-law slopes are steeper at higher baselines compared to the shorter ones, for both field-aligned and perpendicular baselines (Figure 6). We find in Figure 6 that the diffractive scale ($r_{\text{diff}}$) remains much larger for baselines aligned with the magnetic field than for perpendicular ones. Hence, we conclude that the dominant ionospheric structures are elongated along the perpendicular direction (as observed in Figure 5 and 6).
@@ 4 Summary and Conclusions
The analysis presented here uses uGMRT Band 4 data (575–725 MHz), closely following the phase structure function methodology of Mevius et al. (2016), but applied to a distinct frequency regime and instrument. This work extends phase structure function studies to a higher frequency range compared to LOFAR’s 150 MHz observations.
The differing observing frequency range has important implications. Ionospheric phase fluctuations scale inversely with frequency, resulting in smaller phase RMS and often smaller outer scales at higher frequencies for identical ionospheric conditions. Consequently, phase stability tends to improve swiftly as frequency increases, affecting the required calibration intervals and the characteristics of direction-dependent errors.
@@
To further assess anisotropy in the phase fluctuations, we compared the phase structure function along baselines that are approximately aligned with the Earth’s magnetic field (0–20) and those nearly perpendicular to it (70–90). The diffractive scale ($r_{\rm diff}$) is much larger for baselines aligned with the magnetic field than for those perpendicular to it (Figure 5). This means that the ionospheric phase fluctuations are smaller and less variable along the field direction compared to across it. This suggests that the dominant ionospheric structures are elongated along the perpendicular direction - that is, they have shorter correlation lengths (resulting in more rapid phase changes) perpendicular to the field, and are more coherent (showing less variation) along the field direction.
@@
The uGMRT’s Band 4 is also particularly relevant for tied-array Very Long Baseline Interferometry (VLBI) operations, especially in synergy with the Square Kilometre Array (SKA). Tied-array modes require phase coherence across the array, limited by atmospheric and ionospheric fluctuations. Measuring phase structure functions in Band 4 directly informs direction-dependent calibration (DDC) solution intervals and spatial scales necessary for maintaining coherence, helping to optimize calibration overheads for SKA-VLBI and uGMRT tied-array operations.
The implications for tied-array VLBI and direction-dependent calibration are well supported by recent studies and technical discussions in the literature. Key works informing this context include ionospheric calibration and beam-forming techniques for tied-array processing (Ord et al., 2019), VLBI calibration frameworks for SKA and uGMRT (Rioja et al., 2018; Rioja and Dodson, 2022; Li et al., 2024)(e.g., Rioja et al. 2018; Li et al. 2024), and ionospheric error modelling algorithms (Intema, H. T. et al., 2009). These references collectively underscore the necessity of robust direction-dependent calibration for maintaining phase coherence in tied-array modes, especially at mid frequencies relevant to uGMRT Band 4 and SKA operations.
This study thus provides novel insight into the phase stability regime at intermediate frequencies, bridging low-frequency LOFAR results and anticipated mid-frequency SKA operations, and contributes important empirical constraints for tied-array DD calibration strategies in future interferometric experiments. It is important to note that this study used only nighttime observations taken during the winter season. Therefore, our results do not account for diurnal or seasonal variations. Overall, our results demonstrate the capability of the uGMRT for ionospheric studies at mid-latitudes close to the geomagnetic equator (Opio et al., 2015).
@@ Appendix A Sub-Band Variations in Ionospheric Phase Structure Function (Band-4)
In this appendix, we present additional structure function analyses at different sub-bands within the uGMRT Band-4 frequency range to assess the frequency dependence of ionospheric phase fluctuations. Figures 7, 8, and 9 illustrate the results for two representative frequencies: 612.5 MHz and 712.5 MHz. These include the one-dimensional phase structure function with power-law fits, two-dimensional anisotropic structure functions highlighting directional differences in turbulence, and binned phase variance plots as a function of baseline orientation relative to the projected Earth’s magnetic field. Earth rotation causes the projected baseline length and orientation of a fixed antenna pair to evolve during the scan on 3C48. To capture the ionospheric variation within shorter time scales we also repeated our analysis within $\sim$ 2 hour time window (de Gasperin et al. (2018), Figure 10). We find the power law slope is mostly similar, even though the diffractive length scale changes mostly from low to higher values, which suggest a quieter ionosphere as our observation progresses. A similar trend also can be seen in Figure 1, where the variation in $\delta{\rm TEC}$ values are more rapid towards the start of the observation. Together, these plots provide insights into how the ionospheric irregularities and anisotropic characteristics evolve across Band-4 frequencies.
@@
Figure 10: The phase structure function is divided into five time chunks. The initial part of the observation is visibly more affected by ionospheric disturbances, as reflected by the lower $\rm r_{diff}$} values. This trend is also observed in differential TEC values in Figure 1.